OpenStudy (lolbunny):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 2−x and y = 4x + 3 intersect are the solutions of the equation 2−x = 4x + 3. (4 points) Part B: Make tables to find the solution to 2−x = 4x + 3. Take the integer values of x only between −3 and 3. (4 points) Part C: How can you solve the equation 2−x = 4x + 3 graphically? (2 points)

OpenStudy (lolbunny):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 2−^x and y = 4^x + ^3 intersect are the solutions of the equation 2−^x = 4^x + ^3. (4 points) Part B: Make tables to find the solution to 2−^x = 4^x + ^3. Take the integer values of x only between −3 and 3. (4 points) Part C: How can you solve the equation 2−^x = 4^x + ^3 graphically? (2 points)

OpenStudy (lolbunny):

Read second, not first. Forgot to put exponents